[Brahms-l] Shift report 20060618 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 18 2006 - 08:01:05 EDT
Shifter leader was cn
Shifters were nt


Eight hours of beam in this shift 
except for 30~60 minute inturreptions 
for polarization measurement and changing angles
Runs 16210 - 16218 
Run 16219 going on 
No detector problems so far

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 18 08:40:07 2006

This archive was generated by hypermail 2.1.8 : Sun Jun 18 2006 - 08:40:23 EDT